### v2

Abstract
Airdrops are used by blockchain applications and protocols to attract an initial user base, and to grow the user base over time. In the case of many airdrops, tokens are distributed to select users as a “reward” for interacting with the underlying protocol, with a long-term goal of creating a loyal community that will generate genuine economic activity well after the airdrop. Although airdrops are widely used by the blockchain industry, a proper understanding of the factors contributing to an airdrop’s success is generally lacking.
In this work, we outline the design space for airdrops, and specify a reasonable list of outcomes that an airdrop should ideally result in. We then analyze on-chain data from several larger-scale airdrops to empirically evaluate the success of previous airdrops, with respect to our desiderata. In our analysis, we demonstrate that airdrop farmers frequently dispose of the lion’s share of airdrops proceeds via exchanges. Our analysis is followed by an overview of common pitfalls that common airdrop designs lend themselves to, which are then used to suggest concrete guidelines for better airdrops.
keywords:
Airdrop design, Airdrop farmer, Web3, Layer 2
1 Introduction
Blockchain protocols often design reward programs to attract new customers and foster the loyalty of existing users. In recent years, the practice of distributing platform-minted tokens, known as “airdropping,” has gained widespread popularity [53]. For example, in 2023 alone, 4.56 billion US dollars worth of tokens were airdropped to users across various protocols [9]. Although the use of airdrops is prevalent in the blockchain world, we present preliminary findings showing that there is no significant correlation between performing an airdrop and a platform’s popularity relative to available alternatives. Intuitively, this is not optimal and may result in the loss of funds that can be used to improve the Quality-of-Service (QoS)offered by the platform.
Although the fundamental concept of airdrops is relatively straightforward, the design space for such reward schemes is expansive and may vary depending on the characteristics of the platform implementing the airdrop. For instance, certain airdrop mechanisms focus on power users, granting them a substantial amount of rewards with the expectation that these users will stimulate valuable economic activity, subsequently attracting additional users. However, a potential drawback to this approach arises when the distributed tokens empower users to propose changes to a protocol through decentralized governance. It allows users to suggest or vote on changes within a platform, typically employing a one-token-one-vote strategy, where a single user may possess more than one voting token. This approach introduces the risk of voting power concentration, where a minority of users hold the majority of the decision-making voting power [56, 22, 68].
To understand why previous airdrops did not always achieve their goals and to quantify their success, we first articulate a set of reasonable desired outcomes for airdrops. Armed with a basic desiderata, we then go over previous airdrops, see how they measure up, and uncover some interesting insights in the process. Specifically, we analyze the data pertaining to five popular airdrops (ENS [19], dYdX [39], 1inch [10], Arbitrum [41], and Uniswap [1]) and a fake airdrop conducted by a sybil farmer (Gemstone [35, 50]). We conclude that most of the proceeds (as high as 95%) have been sold via exchanges shortly after the airdrop, implying that the airdrops fall short of their intended purpose and benefit mostly airdrop farmers, highly-skilled users who employ sophisticated tactics to increase the share of airdropped tokens that they receive. Furthermore, we describe the common challenges faced by airdrops in the past. Given that the blockchain ecosystem is still quite young, it is expected that the theory and practice of blockchain airdrops are still in its infancy and thus that previous airdrops may not have been entirely successful in the long run. Finally, we use our analysis to suggest potential ways to create better airdrop mechanisms that are more fair to honest users.
1.1 Research Questions
In this section, we address three critical research questions central to our study.
[…]
We evaluate the effectiveness of blockchain protocols in converting short-term user incentives, such as airdrops, into long-term user engagement. The primary goal of conducting an airdrop is to attract users in the short-term and encourage them to become long-term users. Our investigation aims to measure how well these protocols achieve this objective.
Contrary to the assumption that airdrops are the primary driver of long-term user base growth, our research reveals otherwise. Through an exploration of past airdrops conducted, we demonstrate that airdrops may not be the sole determinant of sustained user retention. For example, findings from a recent report by Messari [8] indicate that users join a platform when short-term incentives are available, such as airdrops. However, these protocols struggle to retain users in the long-term, as users often shift their attention to protocols offering better incentives as soon as the initial rewards expire. This behavior implies that protocols successfully onboard users during incentive periods, like airdrops, but witness a drop in user engagement once the incentives come to an end.
Additionally, protocols not conducting airdrops also manage to onboard users due to an expectation of potential future airdrops. Our findings align with this observation, as our analysis indicates that, even after Arbitrum conducted an airdrop, other protocols attracted user attention and increased their total value locked (TVL). This suggests that airdrops alone may not be sufficient to ensure sustained user retention of the platform performing the airdrop, and other factors play crucial roles in long-term platform success.
[…]
Since protocols face challenges in achieving sustained user retention through airdrops alone, an alternative strategy involves incentivizing developers and innovators to build new applications, such as decentralized exchanges (DEXs)and non-fungible token (NFT). This approach aims to foster a dynamic ecosystem that benefits all users on the platform. To enhance user experience and maintain user loyalty, a focus on developing more, potentially exclusive, decentralized applications (dApps) is crucial. Another interesting approach is to provide additional incentives, such as discounts on fees for future platform interactions. For example, in the context of a rollup, users who receive airdropped tokens could redeem them for discounted fees, encouraging long-term engagement. Also, implementing multi-round airdrops allows protocols to learn from previous rounds, enabling iterative improvements to address previously identified issues [2]. Community engagement also plays a vital role in understanding the needs and preferences of users.
In line with these considerations, Blast has introduced an innovative airdrop strategy centered around user contributions to their protocol [55]. This strategy involves awarding points to users as a quantitative measure of their contribution. Points are granted for actions such as successfully bringing in new participants to the blockchain protocol, similar to a referral system, or for users who bridge (send) their tokens from one chain to Blast.
Furthermore, user engagement in blockchain protocols often hinges on short-term incentives. Users tend to join the platform when these incentives, such as airdrops, are available. However, a notable trend is the propensity of users to migrate opportunistically from one blockchain to another, following the trail of incentives. Consequently, a platform experiences an influx of users during incentive periods, only to witness a decline when these incentives cease.
[…]
A crucial aspect to consider involves airdrop farmers. These are sophisticated users who employ intricate tactics to maximize their chances of receiving more tokens per airdrop. They often create numerous wallet addresses controlled by a single entity, a practice known as Sybil [37, 64]. Detecting Sybils can be challenging [21, 51, 36, 73, 59], especially when these accounts execute automatic and randomized actions (e.g., token transfers, bridging, providing liquidity, exchanging tokens, NFTs, etc.) on the blockchain without linking their wallets. Furthermore, the accessibility of airdrop farming software facilitates the automated execution of such actions [70, 29, 12, 15]. Given its intricacy, current methods for identifying Sybils predominantly rely on graph network analysis [21, 51] and machine learning techniques [36, 73, 59].
Another common challenge involves airdropping governance tokens. These tokens empower holders to propose or vote on protocol changes. Distributing these tokens can be risky when users do not prioritize the protocol’s best interests. Additionally, it poses the danger of granting excessive power to certain users, as these protocols typically use a one-token-one-vote mechanism [56, 22, 68]. While acknowledging the challenges highlighted earlier, we believe that these challenges can serve as valuable lessons for protocols venturing into the realm of airdrops, offering insights into the potential path to success.
Furthermore, it is essential to uphold a positive reputation within the community by avoiding any occurrences of insider trading [5, 42]. Insider trading refers to individuals within an organization using privileged information to gain a financial edge over other protocol users. This practice is a common violation of securities laws in traditional financial markets [6], and engaging in it can lead to negative reactions from the blockchain community [69, 61, 23].
Finally, the points discussed above raise questions about whether protocols are effectively achieving their goals post their airdrop campaigns.
1.2 Summary of our Contributions
We summarize our contributions as follows.
[…]
2 Airdrop Desiderata
Airdrops are powerful tools for promoting protocols and acquiring users, attracting newcomers, and incentivizing existing users to engage with the blockchain protocol or application. They are widely utilized for these purposes, with numerous examples documented in the literature (see Table 1). Any protocol capable of creating tokens can utilize airdrops to distribute them to users. For example, blockchain rollup solutions like Arbitrum and Optimism, as well as DeFi applications such as Uniswap, 1inch, dYdX, and ENS have all employed airdrops.
Airdrops can manifest in various forms, with single-round and multi-round distributions being the most prevalent [2]. In a single-round airdrop, tokens are distributed to users all at once, whereas multi-round airdrops involve distributing tokens over multiple rounds each following different strategies. This approach allows leveraging insights from previous rounds to address challenges encountered, such as mitigating potential sybil (i.e., multiple accounts controlled by a single entity) by observing past user behavior patterns [2]. The choice between single-round and multi-round airdrops depends on the protocol’s objectives and community dynamics. Community engagement is a critical aspect of airdrop development; it should be carefully considered by protocol developers.
Moreover, the timing of an airdrop significantly impacts the number of eligible users receiving tokens. Established protocols opting for delayed airdrops may have a larger user base compared to newer projects executing airdrops sooner. This difference in user base size introduces an added layer of complexity when performing sybil detection, as there are more accounts to evaluate and potentially safeguard against. If not performed correctly, this could impact community happiness. In such instances, the protocol may inadvertently reward accounts linked to industrial farming, which could be viewed negatively by the community. However, addressing industrial farming poses significant challenges and complexities. To mitigate this issue, LayerZero Labs implemented a self-reporting mechamism for sybils. Under this system, sybils have the option to self-report and receive 15% of their designated token allocation. Failure to self-report may result in them not receiving tokens if they are identified by the company or its affiliates [46].
Table 1: List of popular airdrops we analyze in this paper.
[…]
dYdX Sept. 2021 Still active Ethereum Continuous DEX
1inch Dec. 2020 Still active Ethereum Continuous DEX
[…]
Arbitrum Mar. 2023 Sept. 2023 Arbitrum One Single Single
Uniswap Sept. 2020 Still active Ethereum Continuous DEX
[…]
Next, we break down airdrop’s high-level goal to bootstrap a community of users into multiple sub-goals. These sub-goals are not necessarily orthogonal, and furthermore, there may be others that are worthwhile to pursue; we focus on them as they serve to illustrate interesting issues faced by common airdrop mechanisms.
Attract users in the short-term.  Historically, airdrops have been employed by new blockchain protocols to establish an initial user base and to provide an initial boost of liquidity to the underlying chain and its protocols. Decentralized platforms, in particular, tend to become more useful and attractive to users as the level of economic activity increases.
Foster long-term involvement.  Bootstrapping an initial base of users is important, but not enough for sustaining a high degree of economic activity over time. Ideally, users should become regular, everyday users. This can be achieved by issuing rewards that can only be used within the blockchain protocol or application, similarly to frequent-flyer points. For example, in layer-2 blockchains (L2s) [67], one may award discounts on future transactions. Other measures that may help are having multiple airdrops, and rewarding users who perform “quests” that expose them to all that the blockchain has to offer. For example, Linea’s airdrop quests gave users an in-depth tour of their various features and use-cases [44].
Target value-creating users.  An airdrop should ideally focus on users who can contribute the most to the long-term sustainability of their platform. In the context of platforms that rely on user-based liquidity provisioning, this may refer to users who provide the most liquidity to lending pools [78] and DEXs [77], or across a variety of tokens. In rollups, one may consider so-called “creators” who deploy popular and useful contracts as particularly valuable. Such users provide additional use cases for a platform, and thus serve to attract other users.
3 Exchange Aftermarket Analysis
The motivation for our quantitative analysis in this paper comes from the following observation: airdrop recipients quickly sell tokens and jump ship, which obviously defeats the point of airdrops in the first place. Analyses performed on DEXairdrops show that airdrop recipients sometimes sell all of their tokens shortly after receiving them. For example, 75% of the recipients of Uniswap’s airdrop dumped all of their tokens within a week of receiving them [38], and 61% have sold their tokens shortly after the drop in the case of ParaSwap [21].
In both instances, the majority of recipients ceased using the relevant blockchain protocols within a few months. This pattern suggests that the airdrop was ineffective in keeping recipients engaged, or that many recipients were likely sybil accounts. Additionally, a rapid sell-off could unsettle the market, particularly if seen as an indication of dwindling confidence in the protocol’s future prospects. Here, we examine data associated with six airdrops (ENS, dYdX, 1inch, Arbitrum, and Uniswap, including Gemstone) presented in Table 1, which were gathered from archive nodes of Ethereum and zkSync Era.To identify exchanges, we use a list of 484 exchange account addresses on Ethereum’s mainnet obtained from Dune using the query: select * from addresses_ethereum.dex and from Etherscan [20].
Table 2: General airdrop distribution statistics.
[…]
supply distributed exchanges addresses
[…]
dYdX $1\times 10^{9}$ $1.77\times 10^{8}$ 35.45% $42682$
1inch $1.5\times 10^{9}$ $83.2\times 10^{6}$ 54.05% $43402$
Gemstone $50.6\times 10^{15}$ $50.4\times 10^{15}$ 95.01% $20014$
[…]
Table 3: Tokens received per address.
[…]
dYdX $2325.85$ $91336.75$ $392.56$ 00 $24.7\times 10^{6}$
1inch $1917.86$ $50164.49$ $636.48$ 00 $9.75\times 10^{6}$
Gemstone $2.31\times 10^{12}$ $6.93\times 10^{11}$ $2.51\times 10^{12}$ 0 $2.64\times 10^{12}$
Arbitrum $1874.02$ $1566.92$ $1250$ $625$ $10250$
Uniswap $620.13$ $9347.91$ $400$ $400$ $2.1\times 10^{6}$
[…]
3.1 Tokens Distributed Over Time
Table 2 presents the number of airdrops distributed and subsequently sold on exchanges. Additionally, Table 3 details the distribution of airdrops per user. These tables provide quantitative insights into how airdropped tokens are disseminated across the user base. Our analysis reveals a significant number of recipients for each airdrop, suggesting potential involvement of airdrop farmers (see column 5 in Table 2). Furthermore, the data indicate that tokens were frequently traded on exchanges. This is supported by the observation that the initial transfer made by the addresses post-airdrop was often a sale on an exchange (see column 4 in Table 2). Notably, the case of Gemstone is particularly striking, with 95% of its tokens being sold on exchanges. In this particular Gemstone case, a non-open sourced DEX created by the airdrop farmer [35].
Furthermore, Gemstone’s airdrop stands out significantly in terms of the number of tokens distributed compared to other airdrops. This substantial distribution also results in a notably higher median number of airdrops per recipient when compared to the other airdrops as shown in Table 3. Remarkably, Gemstone allocated the lion’s share, specifically 99.53%, of their total token supply during their airdrop campaign. It is important to highlight that Gemstone was executed primarily as a sybil attack rather than a legitimate airdrop.
Figure 1: Tokens claimed per day for each of the Airdrops analyzed.
Among all the airdrops analyzed in this study, the Arbitrum airdrop exhibited the fastest rate of token claims by users. Figure 1 illustrates the distribution of accounts claiming their tokens each day. Arbitrum conducted a substantial airdrop, distributing $1162166000$ ARB tokens among $625143$ selected accounts. Notably, $583137$ (93.28%) of these accounts successfully claimed 94.03% of the allocated ARB tokens. Remarkably, 72.45% of the accounts completed their token claims on the first day, with an additional 14.41% doing so on the following day. This cumulative figure amounts to almost 87% of accounts claiming the Arbitrum airdrop within one day of its release, indicating high user engagement and prompt action among the majority of participants.
Table 4: Airdrop recipients’ interaction with exchanges.
[…]
Total # sent to exchanges % sent to exchanges
[…]
dYdX $42682$ $41320$ 96.81%
1inch $43402$ $41913$ 96.57%
Gemstone $20014$ $20000$ 99.93%
Arbitrum $583137$ $555040$ 95.18%
[…]
Table 5: Distribution of hops for airdrop recipients connecting with exchanges.
[…]
Mean Std. Median Min. Max.
[…]
dYdX 2.25 0.59 2 1 18
1inch 1.85 0.64 2 1 7
Gemstone 1 1 1 1 1
Arbitrum 2.71 $0.79$ $3$ $1$ $32$
Uniswap 2.11 $0.43$ 2 1 9
[…]
(a) Over the number of post-airdrop blocks.
(b) Over the number of post-airdrop days.
Figure 2: Time for airdrop tokens to be transferred to exchanges.
Users typically interact with exchanges to convert one token for another and to sell them. To assess how often users receiving an airdrop sell these tokens on exchanges for profit, we analyzed user interactions with exchanges post-airdrop. Table 4 indicates that a significant proportion of airdrop recipients engage with exchanges to trade their tokens. This engagement ranges from a minimum of 83.79% in the case of ENS to a maximum of 99.93% for Gemstone.
Additionally, Table 5 presents the shortest path from each recipient address of an airdrop to any exchange address in our dataset. Our findings reveal that the transfer of tokens to exchanges generally involves only a few steps, suggesting that airdrop recipients do not make significant efforts to conceal their activities. Gemstone is a notable outlier, with all tokens being sold within a single hop. Surprisingly, most accounts managed to reach an exchange with a relatively low number of intermediary hops (between 2 and 3 hops). This observation underscores the pivotal role that exchanges play within the cryptocurrency ecosystem.
(a) ENS.
(b) dYdX.
(c) Gemstone.
(d) 1inch.
(e) Arbitrum.
(f) Uniswap.
Figure 3: Network structure for selling airdropped tokens in 6 airdrops.
Figure 2 shows the cumulative distribution functions (CDFs) demonstrating how quickly users interact with an exchange right after receiving tokens from an airdrop. As shown in Figure 2(a), the majority of accounts exchanged their tokens within approximately one million blocks. The number of blocks for Gemstone is significantly higher than others due to a delay introduced by airdrop farmers. Given the difference in block generation times between Ethereum (with a new block created approximately every 15 seconds) and zkSync, we normalized the blocks in terms of days. According to Figure 2(b), within a single day, 64.67% of the accounts interacted with an exchange. In comparison, ENS had an interaction rate of 45.86%, while dYdX stood at 37.41%. This rapid exchange behavior does not align with one of the primary objectives of a protocol conducting an airdrop, which is to promote sustained user engagement, as discussed in Section 2. The prompt exchange of tokens suggests that users might be leaving a protocol soon after receiving their airdrops.
3.2 Token Distribution Graphs
To better understand the structure of the transfers that happened after each address received the airdrop tokens, we analyzed the transfer network. Denoted as G(V, E), every node (V) represents an address, and an edge (E) is established when tokens are transferred from one address to another. Specifically, the ENS network consists of $184585$ nodes and $608462$ edges, the dYdX network comprises $112853$ nodes and $406027$ edges, the Gemstone network encompasses $20003$ nodes and $240113$ edges, and the 1inch network features $308329$ nodes and $1400913$ edges. Lastly, Arbitrum accounts for $921650$ nodes and $27438608$ edges while Uniswap contains $1180830$ nodes and $3762613$ edges.
To make these graphs more visually readable, we limited the number of hops in the data to one hop from any address that received the protocol airdrop the first few hours after the airdrop and plot their largest component in Figure 3. We manually labeled the higher in-degree nodes using labels provided by Etherscan, a popular blockchain explorer [20]. We show that for all protocols but Gemstone, the decentralized exchange that received the majority of transfers (measured in terms of in-degree) from airdrop recipients is Uniswap, followed by SushiSwap.
[…]
(a) Absolute number.
(b) Relative daily active addresses.
Figure 4: Unique daily active addresses per protocol: (a) Absolute number; (b) relative to the average for the 50 days before Arbitrum’s airdrop.
Figure 5: Ratio between the number of unique daily active addresses in Arbitrum and Optimism.
(a) Daily median fees.
(b) Daily average fees.
(c) Average fees spent per unique address per day.
Figure 6: Daily transaction fees in USD. (a) Median; (b) Average; and (c) Per unique address.
4 Measuring the Airdrop Lift
Empirical evidence suggests that some airdrops excel in attracting users in the short-term, at least superficially [21]. Although there is some preliminary data showing that airdrops fail in achieving the other goals, there is a dearth of solid research on the topic [30, 21]. In this section, we examine this through the lens of some relevant metrics applied to the Arbitrum airdrop. We rely on data source from Growthepie [34]. This platform supplies various metrics, including daily transaction count, daily active addresses, median transaction cost, TVL, fees paid by users, and stablecoin market cap, for several key protocol such as zkSync Era, OP Mainnet, Arbitrum, Immutable X, and Polygon zkEVM. To enhance clarity, all these metrics were subjected to a 7-day moving average smoothing process.
4.1 Unique Active Addresses
Figure 4 examines the number of unique addresses that were active on each protocol using two different approaches. The Figure 4(a) depicts the absolute number of addresses. The Figure 4(b) shows the relative amount of unique daily active addresses for each protocol, when compared to the same protocol’s average number of unique daily active addresses for the period between February 1st 2023 and March 22nd 2023 [66]. This latter date is particularly interesting, as a day later Arbitrum distributed its airdrop to network users [7].
{observation}
others outperformed Arbitrum, without an airdrop. The absolute number of addresses has increased for Arbitrum post-airdrop, and remained at least 50% higher than its pre-airdrop level in the four months since then. However, by looking at other protocols, it becomes apparent that a similar result can be attained without an airdrop. This is demonstrated by Optimism’s trajectory in Figure 4(b), which exhibits an even larger increase in the number of addresses shortly after the beginning of May, perhaps in anticipation of Optimism’s Bedrock launch in the beginning of June 2023 [62]. This is further exemplified by zkSync Era overtaking Arbitrum’s in Figure 4(a), two months after the airdrop.
{observation}
fees may explain the closing gap between Arbitrum and Optimism. Arbitrum leads over Optimism with respect to the number of unique daily active addresses. We turn our focus to Figure 5, which depicts Arbitrum’s lead over Optimism with respect to the number of unique daily active addresses. While in the 50 days before the airdrop Arbitrum has enjoyed an average of 2.6 times more unique daily active addresses than Optimism, this gap has recently started to close: the average over the last 50 days in our dataset is considerably lower, and is equal to 1.83 times more. Optimism’s slightly lower daily median transaction fees since the beginning of June may partly explain their catching up with Arbitrum, as depicted in Figure 6(a). We note that with respect to the average fee, the gap between the protocols is even smaller over the same period as shown in Figure 6(b).
{observation}
the number of unique addresses oscillates over time. Some of the previous observations exhibit oscillatory behavior, in particular in Figure 4(b) depicting the relative amount of unique addresses: rising, reaching a peak, and then subsequently dropping, all in rapid succession. In particular, the phases of Optimism and Arbitrum are inverted regarding the relative amount of unique addresses. One may attempt to explain this by reasoning that when a peak is reached in one of the protocols, fees should increase due to the load, thus prompting cost-conscious users to turn to a different protocols. But, when considering the median transaction fees paid per day (per Figure 6(a)), it is apparent that fees do not have corresponding peaks and dips. In fact, Arbitrum consistently had lower median fees in mid-May.
[…]
4.2 Transaction-related Metrics
A natural next step is to examine transaction-related metrics. In particular, these may serve as a proxy for “real” economic activity which users are willing to pay for, as sending a transaction requires paying some minimal fee [32], when ignoring retroactive rebates or the possibility of blockchain protocols operating at a loss [63].
{observation}
[…]
Some notable differences are that with respect to transaction counts, the gap between Arbitrum and Optimism has nearly closed entirely by the end of July. Furthermore, Figure 7(b) shows that the daily transaction count for Immutable X has almost halved since Arbitrum’s airdrop. Contrast this with the corresponding figure for unique addresses, which shows that on average that metric for Immutable X did not change much, hinting that users have become less engaged with the protocol.
(a) Absolute daily transactions.
(b) Relative daily transactions.
Figure 7: Daily number of transactions per protocol. (a) Absolute number; (b) relative to the average for the 50 days before Arbitrum’s airdrop.
{observation}
Arbitrum’s per address transaction count dropped since the airdrop. To better understand how Arbitrum’s airdrop affected user engagement in other protocols, we now present the Figure 8. Figure 8(b), which depicts the relative average daily transaction count per unique address. The figure shows that since the drop, Arbitrum’s per-user transaction count has fallen to less than 75% of its pre-airdrop level.
(a) Average number.
(b) Relative to average before Arbitrum’s airdrop.
Figure 8: Daily transactions per address.
[…]
{observation}
since June, average transaction fees are similar for all protocols. Conceptually, good metrics should measure a user’s amount of skin in the game. A possible proxy for that is the amount of funds the user is willing to expend for interacting with the protocol. When considering the relative metrics, such as the average fee per transaction or the average fees unique address, one can observe that since June, all protocols are quite similar in both regards. Furthermore, when comparing the average fee per transaction to the median one (per Figures 6(a) and 6(b)), it is apparent that the median fee may be slightly more useful when trying to discern causes for shifts in user behavior.
Another interesting metric is the relative average fees per address, compared to the average in the 50 days preceding Arbitrum’s airdrop, shown in Figure 9. One can observe that Arbitrum’s user engagement was not significantly influenced by the airdrop, and usually behaves similarly to other protocols.
Figure 9: Average fees spent per unique address per day, relative to average before Arbitrum’s airdrop.
(a) Total daily transaction fees (USD).
(b) Relative to average before Arbitrum’s airdrop.
Figure 10: Total fees before and after the Arbitrum airdrop.
Figure 11: Ratio between the daily fees spent in Arbitrum and Optimism.
{observation}
Arbitrum’s total daily fees spiked around the time of the airdrop. Figure 10 shows that Arbitrum’s airdrop did not confer a substantial advantage over its competitors when it comes to its income stream from transaction fees. Although the total amount of fees it earned on the day of the drop reached a peak, this was momentary.
Indeed, per Figure 11, in the 50 days before the airdrop Arbitrum earned on average 1.96 times more from transaction fees per day, while in the last 50 days of our data set this has shrunk to 1.74.
(a) Daily TVL, in USD.
(b) Relative to the average prior Arbitrum’s airdrop.
Figure 12: Total Value Locked (TVL) growth: (a) Daily TVL in USD; and (b) relative to the average for the 50 days before Arbitrum’s airdrop.
4.3 Total Value Locked
The Total Value Locked (TVL) of a protocol is a metric that measure the combined value of all assets stored in that protocol.
{observation}
[…]
5 Common Airdrop Design Challenges
Airdrops share similarities with traditional loyalty programs, such as newcomer bonuses offered to new customers by banks and credit card companies. Furthermore, they also share some common design challenges. The blockchain setting and the mechanics adopted by most airdrops may exacerbate these pitfalls, or introduce new ones.
5.1 Airdrop Farmers
Sophisticated users known as airdrop farmers employ elaborate tactics to maximize the amounts of airdrop tokens they receive. Blockchain protocols have adopted various disproportionate measures to curb the possibility of airdrop farmers gaming the system. For example, a popular approach is to limit the amount of rewards that a single user can receive.
Thus, protocols resort to Proof-of-Humanity services, such as Gitcoin passport [33]. These usually assign a numeric score to users based on a certain metric, where a higher score means that there is more certainty that the corresponding user is indeed human. Gitcoin passport’s metric is based on a list of tasks, such as connecting a user’s social media accounts, or being in possession of some amount of ETH [33]. Such methods can be augmented by analyzing on-chain data to detect sybils and exclude them from the drop [7], but this may lead to false negatives [76]. There are some recent examples of protocols like Celestia considering using sybil resistance solutions [47].
There are other farming mitigation techniques, such as requiring users to perform tasks, ranging from sending a certain type of transactions [65], to sharing posts on social media [45]. These tasks sometimes appear arbitrary, leading to frustration among users. The problem with the aforementioned methods is that they are easy to automate [40] and cheaply fooled, even more-so when protocols offer retroactive rebates on transaction fees [63]. Furthermore, the reliance of many protocols on a limited number of Proof-of-Humanity services means that a one-time investment by an airdrop farmer can result in large profits from multiple airdrops, even with biometric identification, full sybil resistance is not guaranteed [72]. A different approach used by protocols is to announce airdrops that retroactively reward users that were active before the announcement. Nevertheless, farmers can prepare in advance by interacting with these protocols even when no airdrop has officially been announced [14], as was the case in dYdX’s airdrop [18].
The phenomena of reward farming is not limited to crypto-related airdrops, and similar ones can be found in “traditional” loyalty programs. In particular, the practice of credit card churning, where users apply for credit cards for the sole purpose of receiving the rewards bestowed upon newcomers, only to cancel the cards once the rewards were received. Given that similar problems exist even in traditional contexts, where users can be easily identified and penalized, it seems that reward farming cannot be solved easily.
5.2 Decentralized Governance Threats
Some protocols airdrop governance tokens, as a means to decentralize their governance process. However, distributing governance tokens may be dangerous. Such tokens allow holders to vote in the protocols’ underlying governance mechanisms, but usually can be exchanged for other tokens, and thus may potentially have monetary value [17].
There is some empirical evidence that suggests that airdrops of governance tokens may be more successful than airdrops of non-governance tokens. In a recent analysis of the market capitalization of crypto tokens performed by Makridis et al. [53], the authors found that if an airdrop is of a governance token, the token’s market capitalization growth rate can increase by $14.99\%$ relative to governance tokens that were not airdropped. However, the authors acknowledge that this effect is statistically insignificant when using common benchmarks.
Unfortunately, airdropping governance tokens runs the risk of giving too much power to some users, leading to an unfair distribution of decision-making power in the system [56, 22, 68]. Also, to users that do not have the protocol’s best interests in mind, and may even act to the determent of the protocol.
5.3 Insider Trading
It is a problematic occurrence within organizations where individuals exploit privileged information to gain a financial advantage over other users of the protocol. This practice is widely recognized as a violation of securities laws in traditional financial markets [6], and engaging in such activities can lead to negative responses from the blockchain community [69, 61, 23].
When someone within a protocol utilizes privileged information to enhance their profits during an airdrop, it can trigger a backlash from the blockchain community. Insider traders have advanced knowledge of the metrics used to determine eligibility and the amount of an airdrop for each address. This knowledge empowers them to manipulate the system and make a profit. For example, there were claims that a head of growth at AltLayer might have used insider information to make a $200000$ USD profit from their airdrop, although later dismissed as a mere coincidence [69]. However, such incidents can undermine the trust users place in these protocols.
This issue also raises concerns about fairness, as some users have access to superior and more accurate information than others. Identifying these traders is a challenging problem, making it crucial for protocols to provide extensive information to their users. In addition, incentives for blockchain data analytics companies and research groups to conduct post-airdrop data audits can help identify insider traders by analyzing address details and transfer patterns [23]. For this to happen, data availability is key. Therefore, protocols should ensure transparency and encourage thorough analysis to maintain integrity and fairness within the blockchain community.
6 Design Guidelines
The airdrop design challenges outlined earlier, while indeed sobering, may inform blockchain protocols embarking on the airdrop journey, and can shed light on the path to potential success.
6.1 Alternative Incentives may Better Sustain User Engagement
An airdrop’s long-term benefit to a protocol can be tricky to measure, given that the potential benefits may be indirect and hard to quantify, while expenses are partially immediate and irrevocable, e.g., the value of airdropped tokens is transferred from protocols to users and is thereafter no longer available for other uses. Furthermore, some costs and effects cannot be easily foreseen, like those arising from distributing governance tokens to users.
Instead of using airdrops, a community could adopt other measures which have a more predictable relation between costs and benefits. For example, a very natural alternative to airdrops is for the community to vote to programmatically reward power users with discounts on future interactions. When considering L2s, these discounts can be given on transaction fees. Thus, if a user receives this reward, it necessarily has to use the protocol again in the future to benefit from the incentive, thus effectively fostering continued user engagement. Furthermore, this incentive mechanism, although not completely resistant to airdrop farming, does fare better: the discounts have no intrinsic value outside the protocol, and the protocol’s costs go solely towards users actively using the system.
Nevertheless, while fee discounts are more focused towards incentivizing continued interaction within the protocol, they are not without their own challenges. For example, one should carefully design a proper discount mechanism, covering both the eligibility criteria for users, and the height of awarded discounts. Furthermore, it is not immediately apparent that discounts can attract users to a similar degree as the more immediate and tangible rewards offered by airdrops. Another option is to perform multiple airdrops over a relatively long period of time, instead of a single one. Although this is still susceptible to some of the pitfalls of standard one-shot airdrops, it may serve to ensure the long-term engagement of the community, and thus avoid the decline in user adoption that some experience immediately after an airdrop.
A more innovative approach was adopted by Blast, which introduced a point-based reward program [55]. Under this program, users receive rewards based on the points they accumulate through various activities, such as bridging tokens to the protocol (i.e., transferring users’ funds from another protocol) and participating in a referral program where bringing in more users increases their points. Interestingly, Blast received 1.1 billion USD in deposits even before its official launch [43]. This method provides an evaluative metric for users’ contributions to the protocol, structured around a referral-program model.
Furthermore, the incorporation of innovative distribution mechanisms demonstrates as a critical role in mitigating the presence of sybils within whitelisted addresses eligible for an airdrop. For example, Celestia has proposed a distinctive design that employs GitHub commits as a proxy to assess a user’s contribution to the blockchain ecosystem [27]. Nevertheless, a potential concern emerges in regard to the possibility of users or farmers engaging in the generation of counterfeit activities on GitHub, aiming to exploit other protocols’ airdrops adopting a strategy analogous to Celestia’s. Among other metrics, would be to focus on metrics that are resistant to programmable manipulation, adding a layer of difficulty or cost to automated user accounts.
6.2 Target Well-known Reputable Entities
Instead of rewarding pseudonymous users, one can target DAOs that are building relevant apps. In Arbitrum’s airdrop, 1.13% of airdropped tokens were allocated to DAOs [24]. Arbitrum provides additional incentives besides its airdrop for specific populations of interest, such as university students [26] or members of the technical community [25] who wish to take part in research and build tools relevant to the protocol. Another approach was advanced by Optimism, that directed some of its revenue to retroactively fund successful projects, essentially bringing the concept of a startup exit to the blockchain world [28].
6.3 Proactive Oversight and Community Engagement
While the airdrop is ongoing, the protocol’s data should be carefully analyzed to verify that no one is attempting to game the system. A good example was put forth by Linea, whose team uncovered a vulnerability [49] that allowed users to cheat their incentive mechanism. The prompt discovery prevented cheaters from obtaining over one-third [^2] A dashboard is available on Dune at https://dune.com/rusty378/guildxyz-pins. of NFTsoffered as incentives.
A complementary approach is to engage the community and to open communication channels and bug bounties for those who wish to disclose vulnerabilities [71], whether these are actively used or not. For example, a disclosure by an AzukiDAO community member allowed the protocol to mitigate a vulnerability relatively quickly [58]. A protocol’s oversight should extend beyond on-chain data. In particular, social media is rife with scammers who advertise “fake” airdrops for their protocols, where users are promised to receive rewards by connecting their wallets to some innocent-looking site, which actually drains user funds. Even protocols that are not planning an airdrop are targeted by such ploys.
Actively engaging with the community regarding technical matters can prove beneficial, as well. For example, zkSync Era’s NFTdrop was retroactively analyzed by cygaar [16], who found that it could be implemented more cost-effectively. Operating transparently and giving the community visibility into the inner-workings of the protocol can assure that even if technical issues arise, they are met with understanding by users.
6.4 Rewards Should Scale with Costs
The ramifications of Goodhart’s law are readily observable in previous airdrops. For example, some airdrops attempt to incentivize user engagement by publicly announcing that eligibility for rewards is determined according to the number of times users interact with the protocol in the past. Consequently, users can game the system by performing meaningless wash-trading transactions [74], meaning that the eligibility criteria does not serve as an effective measure of a user’s genuine engagement with the protocol.
This is further compounded by the fact that action-based metrics used by airdrops do not usually account for the actual costs incurred by users for each action. For instance, if the number of transactions is used as a metric, one should consider that if transaction fees are low, an airdrop farmer may cheaply perform the required amount of transactions. To address this, adopting a quadratic voting system could deter users from artificially inflating transaction volumes, as their rewards would be proportionate to the square root of their activities [11, 48]. Consequently, the cost to achieve a similar impact would rise, acting as a barrier against manipulation. On the other hand, high fees can eat into the rewards received by users, thereby diminishing the appeal of the airdrop [60]. To remedy these issues, rewards should scale with the costs that are incurred by users.
6.5 Data Availability
Blockchain data is typically publicly accessible but not easily accessible. Ensuring data availability is crucial for monitoring airdrop activities and combating Sybil attacks. While previous research on airdrops and Sybil detection has primarily focused on single blockchains, our study expands this scope to multiple blockchains.
Specifically, we collect data from two major Ethereum rollups, Arbitrum and zkSync Era, and gather and label data from significant airdrops, including a major Sybil attack on zkSync Era (the Gemstone “airdrop”). Since Sybil detection is increasingly challenging and requires data from multiple chains—given that protocols often reward participants from different chains [27, 4]—data availability is essential for analyzing past airdrops and testing new metrics for future applications. However, the rapid growth of blockchain data can make its collection and processing impractical.
Protocols should ensure their data is available for researchers and companies, helping improve monitoring tools and facilitate learning from the data. To support this, we will make our zkSync Era and Ethereum datasets publicly available, enabling others to utilize them for optimizing Sybil detection and airdrop designs.
7 Related Work
The majority of recent papers on airdrops concerned themselves with post-factum analysis of prior airdrops, as well as with suggestions on how to run airdrop campaigns.
Studies.  Makridis et. al. [53] investigate the role of airdrops for DEXgrowth using a sample of 51 exchanges. They find that airdrops of governance tokens, in particular, are positively associated with growth in market capitalization and volume.
Lommers et. al. [52] offer a broad treatment of the different types of airdrops used in practice (plain vanilla, holder-based, past value-based, future value-based). They offer examples of projects like Blur and the pitfalls of their airdrops, along with some advice on airdrop implementation strategies. In particular, they explore how eligibility criteria, signaling, and implementation strategy of the airdrop affect its success, and they provide practical guidance on how to toggle these.
Fan et. al.[21] study the ParaSwap DEXas an example to propose a role taxonomy methodology through a data-driven study to understand the characteristics of community members and the effectiveness of airdrop. They find that users who receive more rewards tend to take positive actions towards the community. They summarize several arbitrage patterns and confirm that the current detection is not sufficient to detect airdrop hunters. They use an unsupervised hierarchical clustering method to capture the behaviors of eligible airdrop addresses after receiving rewards and to divide them into interpretable categories. Other works focused on addressing the problem of Sybil detection, providing valuable insights into identifying airdrop farmers. Commonly adopted practices include graph network analysis [21, 51] and machine learning techniques [36, 79, 73, 13, 59]. The former involves identifying prevalent transfer patterns, such as sequences, radial structures, or a combination of both. The latter utilizes approaches like node embedding and employs universal pre-trained transformer encoders like BERT for effective fraud detection [36] or clustering [73, 59].
Allen [2] examines airdrop design as co-evolutionary process and makes three contributions. First, applying a co-evolutionary lens to airdrops, where participants learn and adapt their strategies (e.g., new criteria, on-chain farming) both within and between web3 ecosystems. Second, nine airdrop cases studies (e.g., Optimism, Arbitrum, Blur) that demonstrate the expanding set of airdrop strategies (e.g., task-based claiming). Third, implications for airdrop design including more feedback loops to facilitate learning (e.g., multi-round airdrops), ongoing innovation in airdrop criteria (e.g., randomness), and a reversion back to simple designs for some projects (because of the costs of airdrop design).
Allen et. al.[3] consider why projects might choose this unusual and costly means of token distribution. It considers a selection of high-profile airdrops as case studies between 2014 and 2022. This is the first comprehensive analysis of the rationales and mechanisms of Web3 token airdrops. They find that two primary rationales for airdrops are marketing (to attract new users and to maintain a community) and decentralization of ownership and control of a project (building community, providing regulatory protection, and enhancing security). While many commentators describe airdrops as primarily for marketing, they claim that marketing is a weak rationale for airdropping tokens and the evidence of successful marketing-driven airdrops lacking.
Technical Aspects of Airdrops  Frowis et. al. [31] point out that efficient transfers to a large number of recipients in the case of airdrops present problems on Ethereum. They find that cost savings of a factor of two are in fact possible, but require specific provisions in the smart contract implementing the token system. Pull-based approaches, which use on-chain interaction with recipients, promise moderate savings for the distributor while imposing a disproportional cost on each recipient. Total costs are broadly linear in the number of recipients, regardless of the technical approach.
Wahby et. al. [75] point out that current airdrops offer no recipient privacy: they leak which recipients have claimed the funds, and this information is easily linked to off-chain identities. In this work, the authors address this issue by defining a private airdrop and describing concrete schemes for widely used user credentials (ECDSA and RSA). Private airdrop for RSA builds upon a new zero-knowledge argument of knowledge of the factorization of a committed secret integer, which may be of independent interest. Their fastest implementation takes 40–180 ms to generate and 3.7–10 ms to verify an RSA private airdrop signature.
8 Conclusions
In this study, we investigated the prevalent issues with airdrops and developed guidelines to improve their impact. Our study analyzed six notable airdrops, uncovering a pattern in which recipients quickly traded their airdropped tokens on exchanges. Specifically, 36.51% of ENS tokens, 35.35% of dYdX tokens, and 53.47% of 1inch tokens were exchanged soon after distribution. Generally, these tokens were sold after only 1 or 2 transactions, with a median of 2 transactions, demonstrating that airdrops do not promote sustained engagement or attract contributors of significant value. Furthermore, our examination of Arbitrum indicated a significant increase in total daily fees during the airdrop event, but a decrease in transaction count per address afterward. In particular, other protocols outperformed Arbitrum in the absence of an airdrop. The average transaction fees across all protocols followed a similar trend since June 2023, suggesting that airdrops may not be the primary driver of user base growth and can impact activity on other chains.
We also acknowledge the common airdrop challenges related to airdrop farmers, governance token distribution, and insider trading. These challenges serve as valuable lessons for blockchain protocols considering airdrop tokens and offer insight into the potential path to success. As a measure to avoid these pitfalls, we propose design guidelines for airdrop campaigns, which are summarized below.
Alternative incentives may better sustain user engagement.  An interesting approach involves offering alternative incentives, such as fee discounts for future interactions within the protocol, to foster long-term engagement. Additionally, implementing multi-round airdrops allows for learning from past rounds, enabling iterative improvements to address previously identified issues. Another approach involves exploring innovative metrics to assess the eligibility of an address for receiving an airdrop. This might also be a promising tactic to address concerns related to airdrop farming who typically tries to game the system by utilizing past airdrop actions. However, this approach may not completely solve the challenge of insider trading.
Target well-known reputable entities.  Prioritizing established and reputable entities, including those building on the protocol, research groups, tech communities, and students, can contribute to sustained engagement. Funding these entities or directing airdrops towards DAOs developing relevant dApps presents an opportunity to attract value-driven users and promote long-term involvement.
Proactive oversight and community engagement.  Critical to the success of any airdrop is proactive oversight and community engagement. Analyzing data from past airdrops, even those on different chains, provides insights into pitfalls and allows for adjustments in future airdrop strategies. Open communication channels with the community help understand their interests and desired protocol functionalities.
Rewards should scale with costs.  Ensure that rewards scale proportionally with associated costs to prevent exploitation by airdrop farmers. Aligning rewards with costs is a strategic measure to mitigate potential gaming of the system.
In conclusion, our paper offers insights into airdrop challenges and proposes practical guidelines for designing effective airdrop campaigns. We hope these suggestions inspire blockchain protocols, researchers, and practitioners. Finally, in our commitment to promoting scientific reproducibility, we plan to share the dataset and scripts utilized in this study [57].